# Distinguishing the origin of eccentric black-hole mergers with gravitational-wave spin measurements

### v1

@@ Abstract
It remains an open question whether the binary black hole mergers observed with gravitational-wave detectors originate from the evolution of isolated massive binary stars or were dynamically driven by perturbations from the environment. Recent evidence for non-zero orbital eccentricity in a handful of events is seen as support for a non-negligible fraction of the population experiencing external driving of the merger. However, it is unclear from which formation channel eccentric binary black-hole mergers would originate: dense star clusters, hierarchical field triples, active galactic nuclei, or wide binaries in the Galaxy could all be culprits. Here, we investigate whether the spin properties of eccentric mergers could be used to break this degeneracy. Using the fact that different formation channels are predicted to either produce eccentric mergers with mutually aligned or randomly oriented black-hole spins, we investigate how many confident detections would be needed in order for the two models to be statistically distinguishable. If a few percent of binary black hole mergers retain measurable eccentricity in the bandwidth of ground-based detectors, we report a $\sim 40\,\%$ chance that we could confidently distinguish both models after the fifth observing run of the LIGO-Virgo-KAGRA detector network, $\sim 80\,\%$ for LIGO A#, and $\sim 98\,\%$ for the Einstein Telescope and Cosmic Explorer.
1 Introduction
@@
Ten years after the first direct detection of gravitational waves (GWs) from merging binary black holes (BBHs) (Abbott et al., 2016), about a hundred merger events have been observed (Abbott et al., 2023a). However, the astrophysical mechanism forming them remains unclear. Multiple formation channels have been proposed, which can be broadly divided into two categories. On the one hand, isolated massive binary stars may develop merging BBHs after they have interacted through a stable mass-transfer episode (van den Heuvel et al., 2017; Inayoshi et al., 2017; Bavera et al., 2021; Marchant et al., 2021; Gallegos-Garcia et al., 2021; Olejak et al., 2021; Stegmann & Antonini, 2021a; van Son et al., 2022), in a common-envelope phase (Postnov & Yungelson, 2014; Belczynski et al., 2016; Eldridge & Maund, 2016; Lipunov et al., 2017), or in a chemically homogeneous evolution (de Mink & Mandel, 2016; Mandel & de Mink, 2016; du Buisson et al., 2020; Riley et al., 2021). On the other hand, BBHs may be driven to merge by some perturbation from the environment, e.g., distant tertiary companions in hierarchical triples (Blaes et al., 2002; Thompson, 2011; Antonini & Perets, 2012; Antognini et al., 2014; Antonini et al., 2014; Prodan et al., 2015; Silsbee & Tremaine, 2017; Antonini et al., 2018; Rodriguez & Antonini, 2018; Grishin et al., 2018; Hoang et al., 2018; Stegmann et al., 2022b, a; Vigna-G\’omez et al., 2025), scatterings in dense star clusters (Samsing et al., 2014; Antonini & Rasio, 2016; Rodriguez et al., 2018a; Zevin et al., 2021a; Chattopadhyay et al., 2023), fly-bys and the tidal field of the host galaxies (Michaely & Perets, 2019; Raveh et al., 2022; Stegmann et al., 2024), or gaseous embeddings in active galactic nuclei (AGNs) (Goldreich et al., 2002; Bellovary et al., 2016; Tagawa et al., 2020b; Grishin et al., 2024).
@@
A powerful smoking-gun signature of the dynamical or otherwise externally perturbed formation of merging binaries is the presence of non-zero orbital eccentricity in a detected GW signal. This is because GW emission leads to extremely efficient angular momentum loss which rapidly circularises the orbit of a binary (Peters, 1964). Thus, any system that retains a substantial residual eccentricity upon entering the sensitive frequency range of a ground-based GW detector must have formed with a much higher eccentricity. A fraction of BBH mergers with such large eccentricities is only expected if they are driven by an environmental perturbation. Conversely, the presence of non-zero eccentricity at the minimum sensitive frequency of ground-based detectors is wholly inconsistent with the expected evolution of isolated massive binary stars (e.g., Belczynski et al., 2002). Due to these considerations, the detection of even a few sources with non-zero eccentricity (or the lack thereof; Zevin et al. 2021a) would act as a “tip of the iceberg” and could be used to put constraints on the fraction of mergers formed from dynamical channels in the entire observed population (cf., Baibhav et al., 2020).
@@
Tentative evidence for compact object mergers with non-zero eccentricity has recently been found in a handful of BBH merger events (e.g., Romero-Shaw et al., 2022; Gupte et al., 2024; Iglesias et al., 2024; de Lluc Planas et al., 2025), though degeneracies with misaligned spin are, at present, hard to eradicate (e.g., Romero-Shaw et al., 2023) (although, see e.g. Divyajyoti et al. 2024a; Morras et al. 2025). If this evidence becomes more robust, it would suggest that a somewhat large fraction of the BBH mergers detected thus far formed through a dynamical formation channel. Here, we consider a scenario in which additional eccentric BBH merger events are confidently detected, and investigate whether measuring their spin properties would further constrain which of the channels that produce eccentric mergers contribute. We focus on the spin properties because, for the subset of eccentric mergers, different formation channels make very distinct predictions. The two dimensionless spin vectors $\boldsymbol{\chi}_{1}$ and $\boldsymbol{\chi}_{2}$ of the black holes and their orbital angular momentum vector $\mathbf{L}$ are either expected to point in three random directions, or the spins are mutually closely aligned ($\boldsymbol{\hat{\chi}}_{1}\approx\boldsymbol{\hat{\chi}}_{2}$) but randomly oriented with respect to $\mathbf{\hat{L}}$, as argued below and illustrated in the right panel of Fig. 1.
@@
Figure 1: Left panel: Distribution of $\chi_{\rm eff}$ for mutually aligned spin directions (red) and random spin directions (teal). Dotted lines indicate equal-mass ($q=1$) and maximally spinning BBHs ($\chi_{1}=\chi_{2}=1$), solid lines assume an extended mass ratio distribution and a spin magnitude distribution that peaks at lower values, which are consistent with previous GW detections (see text). Black markers show three eccentric BBH merger candidates identified by Gupte et al. (2024) (from bottom to top: GW200129, GW190701, and GW200208_22) and orange markers show four candidates identified by Romero-Shaw et al. (2022) (from bottom to top: GW200208_22, GW190521, GW190620, and GW191109). The commonly identified candidate GW200208_22 is highlighted by a filled marker (Romero-Shaw et al., 2025). Right panel: Sketch of uncorrelated (top) and mutually aligned black-hole spins of eccentric mergers (bottom).
@@
    Three random directions of $\boldsymbol{\chi}_{1}$, $\boldsymbol{\chi}_{2}$, and $\mathbf{L}$ are a natural property of eccentric BBH mergers that form in dense star clusters. There, BBHs that retain measurable eccentricity in the bandwidth of current GW detectors merge during resonant three- and four-body encounters that typically involve chaotic exchanges between binary members and scatterings that heavily perturb their orbits (Samsing et al., 2014; Antonini & Rasio, 2016; Rodriguez et al., 2018b; Zevin et al., 2021a; Chattopadhyay et al., 2023). In these chaotic encounters, merging BBHs are randomly assembled without any correlations between the spins and orbital angular momentum. As a result, the directions of $\boldsymbol{\chi}_{1}$, $\boldsymbol{\chi}_{2}$, and $\mathbf{L}$ are independent of one another and effectively randomised.
@@
    In contrast, closely aligned spins $\boldsymbol{\hat{\chi}}_{1}\approx\boldsymbol{\hat{\chi}}_{2}$ and a random direction of $\mathbf{{L}}$ are expected for eccentric BBH mergers that result from the evolution of hierarchical triples. In these systems, the BBHs are formed from massive binary stars whose spins are likely (closely) aligned to one another by stellar processes. The gravitational perturbation from a tertiary companion promotes a subsequent merger by causing large-amplitude von Zeipel-Kozai-Lidov (ZKL) (Zeipel, 1909; Kozai, 1962; Lidov, 1962) oscillations of the BBH eccentricity and the orientation of its orbital angular momentum. BBHs that merge with a residual eccentricity undergo exceptionally strong ZKL oscillations while the spin evolution is highly non-adiabatic, i.e., the orbital angular momentum vector oscillates several orders of magnitude quicker than the spins could follow it through relativistic spin-orbit coupling (Antonini et al., 2018; Rodriguez & Antonini, 2018; Liu & Lai, 2018; Liu et al., 2019). As a result, the eccentric BBH mergers retain the mutually aligned spins they inherited from their binary progenitor stars while the direction of their orbital angular momentum vector is effectively randomised.
@@
In AGN disks, the spins of BBHs are expected to be mutually aligned through gas accretion due to the Bardeen-Petterson effect (Bardeen & Petterson, 1975), in which the alignment of the black-hole spin with a circum-black-hole disk proceeds much faster than the evolution of the spin magnitude (see McKernan et al. 2020; Santini et al. 2023 for related modeling). However, the directions of the orbital angular momentum can be misaligned with the black-hole spin directions depending on the frequency of binary-single interactions. This frequency is influenced by the properties of the AGN disks, especially the merging locations within the AGN disks. Recent models predict frequent interactions with random angular momentum directions (Tagawa et al., 2020a). In these cases, the spin-orbit alignment for the AGN channel is predicted to be similar to that of the three-body channel, and highly eccentric mergers are expected for a fraction of mergers. On the other hand, for the AGN channel, hierarchical mergers are inevitable (Yang et al., 2019; Tagawa et al., 2021; Gerosa & Fishbach, 2021), resulting in high masses and spins for these black holes, which can help distinguish them from triple systems.
In this work, we study how many confident detections of eccentric BBH mergers we would need before the two spin-orbit configurations of (i) three randomly distributed directions of $\boldsymbol{\chi}_{1}$, $\boldsymbol{\chi}_{2}$, and $\mathbf{L}$ and (ii) two randomly distributed directions of $\boldsymbol{\chi}_{1}$ and $\mathbf{L}$ and $\boldsymbol{\hat{\chi}}_{1}\approx\boldsymbol{\hat{\chi}}_{2}$ become statistically distinguishable.
2 Methods
@@
Random directions of $\mathbf{L}$, $\boldsymbol{\chi}_{1}$, and $\boldsymbol{\chi}_{2}$ correspond to uncorrelated distributions $\cos\theta_{1}\sim\mathcal{U}(-1,1)$ and $\cos\theta_{2}\sim\mathcal{U}(-1,1)$. In contrast, random directions of $\mathbf{L}$ but mutually aligned spins mean $\cos\theta_{1}=\cos\theta_{2}\sim\mathcal{U}(-1,1)$. We refer to these models as “uncorrelated” and “mutually aligned” respectively. The difference between the distribution of $\chi_{\rm eff}$ in both models is maximised if we consider the unrealistic but instructive case of equal-mass ($q=1$) and maximally spinning ($\chi_{1}=\chi_{2}=1$) BBHs. This is shown in the left panel of Fig. 1 (dotted lines), which illustrates that the mutually aligned model tends to produce more extreme values near $\chi_{\rm eff}\approx\pm 1$ whereas the uncorrelated model is more centred around zero. Both models resemble one another more closely if we consider more realistic, non-unity mass ratios and spin magnitudes. In Fig. 1, this is shown with solid lines where we draw the mass ratio from a power-law distribution $p(q)\sim q^{\alpha}$ between 0.1 and 1.0 and the spin magnitudes from some Beta distribution $p(\chi_{1,2})\sim\rm Beta(a,b)$. We choose fiducial values $\alpha=1.1$, $a=1.4$, and $b=3.6$ which are consistent with the distributions inferred from the GWTC-3 (Abbott et al., 2023b). As a result, the mutually aligned model retains more pronounced tails of the distribution, whereas the uncorrelated model is more concentrated around $\chi_{\rm eff}=0$. Yet, the two distributions are more similar than in the previous case where $q=\chi_{1}=\chi_{2}=1$ due to a weaker contribution of the second term on the right-hand side of Eq. (1).
@@
For this purpose, we assume a total number of $N_{\rm tot}$ detected BBH mergers, which we set to $2,100$ as a reference for the 5th observing run of LVK (LVK O5) (Broekgaarden et al., 2024), $10^{4}$ for $A\#$ (Abbott et al., 2020), and $10^{5}$ for the Einstein Telescope (Abac et al., 2025) and Cosmic Explorer (ET/CE) (Reitze et al., 2019). We assume that a fraction $0\leq\xi_{\rm ecc}\leq 1$ of them will retain measurable eccentricity in the detector sensitivity window, resulting in $N_{\rm ecc}=N_{\rm tot}\times\xi_{\rm ecc}$ eccentric BBHs (rounded to the next integer). For each detector configuration, we repeat the following steps $10^{4}$ times:
@@
    We produce a synthetic merger population by sampling $N_{\rm ecc}$ eccentric BBHs with mass ratios drawn from a power-law distribution $p(q)\sim q^{\alpha}$ between 0.1 and 1.0 and spin magnitudes independently drawn from some Beta distribution $p(\chi_{1,2})\sim\rm Beta(a,b)$. We assume that the distribution parameters are drawn from $\alpha\sim\mathcal{N}(1.1,1.0)$, $a\sim\mathcal{N}(1.4,0.1)$, and $b\sim\mathcal{N}(3.6,0.1)$. For each of the eccentric BBH mergers we compute $\chi_{\rm eff}$ using Eq. (1) assuming either mutually aligned or uncorrelated spin directions. Lastly, we approximate the posterior distribution of each $\chi_{\rm eff}$ measurement with a truncated normal distribution between $-1$ and $1$, where the adopted standard deviation $\sigma_{\chi_{\rm eff}}$ represents typical measurement uncertainties.
@@
to quantify the confidence to recover the correct model, i.e., the mutually aligned model in the presence of mutually aligned data and the uncorrelated model in the presence of uncorrelated data. In the discussion below, we follow Jeffreys’ scale and consider $\log_{10}\mathcal{B}\gtrsim 1/2$ to indicate substantial evidence for the correct model, $-1/2\lesssim\log_{10}\mathcal{B}\lesssim 1/2$ no particular preference for either model, and $\log_{10}\mathcal{B}\lesssim-1/2$ substantial evidence for the incorrect model.
@@
If not stated otherwise, we assume an eccentric merger fraction of $\xi_{\rm ecc}=0.05$ which agrees with predictions from dynamical channels for current ground-based detectors (e.g., Antonini et al., 2014; Zevin et al., 2021a), and is consistent with the order of magnitude of eccentric candidates in the current population of GW events (Romero-Shaw et al., 2022; Gupte et al., 2024). However, it likely underestimates the fraction of eccentric mergers detectable with the next-generation detectors ET/CE, as they will be sensitive to smaller eccentricities (Dorozsmai et al., 2025). Below, we also study different values for $\xi_{\rm ecc}$. For the measurement uncertainty of $\chi_{\rm eff}$, we assume as a default $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# which is the typical uncertainty of the (quasi-circular) events identified in GWTC-3. We assume that ET/CE will allow measurements that are more precise by one order of magnitude, $\sigma_{\chi_{\rm eff}}=0.02$. Since it is unclear at what precision next-generation detectors will measure ${\chi_{\rm eff}}$ of eccentric events, we also study different values $\sigma_{\chi_{\rm eff}}$.
3 Results

### v2

@@ Abstract
It remains an open question whether the binary black hole mergers observed with gravitational-wave detectors originate from the evolution of isolated massive binary stars or were dynamically driven by perturbations from the environment. Recent evidence for non-zero orbital eccentricity in a handful of events is seen as support for a non-negligible fraction of the population experiencing external driving of the merger. However, it is unclear from which formation channel eccentric binary black-hole mergers would originate: dense star clusters, hierarchical field triples, active galactic nuclei, or wide binaries in the Galaxy could all be culprits. Here, we investigate whether the spin properties of eccentric mergers could be used to break this degeneracy. Using the fact that different formation channels are predicted to either produce eccentric mergers with mutually aligned or randomly oriented black-hole spins, we investigate how many confident detections would be needed in order for the two models to be statistically distinguishable. If a few percent of binary black hole mergers retain measurable eccentricity in the bandwidth of ground-based detectors, we report a $\sim 9\,\%$ chance that we could confidently distinguish both models (Bayes factor $\ln\mathcal{B}>3$) after the fifth observing run of the LIGO-Virgo-KAGRA detector network, $\sim 63\,\%$ for LIGO A#, and $\sim 98\,\%$ for the Einstein Telescope and Cosmic Explorer.
I Introduction
@@
Ten years after the first direct detection of gravitational waves (GWs) from merging binary black holes (BBHs) (Abbott and others, 2016), about a hundred merger events have been observed (Abbott and others, 2023). However, the astrophysical mechanism forming them remains unclear. Multiple formation channels have been proposed, which can be broadly divided into two categories. On the one hand, isolated massive binary stars may develop merging BBHs after they have interacted through a stable mass-transfer episode (van den Heuvel et al., 2017; Inayoshi et al., 2017; Bavera et al., 2021; Marchant et al., 2021; Gallegos-Garcia et al., 2021; Olejak et al., 2021; Stegmann and Antonini, 2021; van Son et al., 2022), in a common-envelope phase (Postnov and Yungelson, 2014; Belczynski et al., 2016; Eldridge and Maund, 2016; Lipunov et al., 2017; Giacobbo and Mapelli, 2018), or in a chemically homogeneous evolution (de Mink and Mandel, 2016; Mandel and de Mink, 2016; du Buisson et al., 2020; Riley et al., 2021). On the other hand, BBHs may be driven to merge by some perturbation from the environment, e.g., distant tertiary companions in hierarchical triples (Blaes et al., 2002; Thompson, 2011; Antonini and Perets, 2012; Antognini et al., 2014; Antonini et al., 2014; Prodan et al., 2015; Silsbee and Tremaine, 2017; Antonini et al., 2018; Rodriguez and Antonini, 2018; Grishin et al., 2018; Hoang et al., 2018; Stegmann et al., 2022b, a; Vigna-Gómez et al., 2025; Stegmann and Klencki, 2025), scatterings in dense star clusters (Samsing et al., 2014; Antonini and Rasio, 2016; Rodriguez et al., 2018; Zevin et al., 2021b; Mapelli et al., 2021; Chattopadhyay et al., 2023; Dall’Amico et al., 2024), fly-bys and the tidal field of the host galaxies (Michaely and Perets, 2019; Raveh et al., 2022; Stegmann et al., 2024), or gaseous embeddings in active galactic nuclei (AGNs) (Goldreich et al., 2002; Bellovary et al., 2016; Fragione et al., 2019; Tagawa et al., 2020b; Grishin et al., 2024; Vaccaro et al., 2024; Gilbaum et al., 2025).
@@
A powerful smoking-gun signature of the dynamical or otherwise externally perturbed formation of merging binaries is the presence of non-zero orbital eccentricity in a detected GW signal. This is because GW emission leads to extremely efficient angular momentum loss which rapidly circularises the orbit of a binary (Peters, 1964). Thus, any system that retains a substantial residual eccentricity upon entering the sensitive frequency range of a ground-based GW detector must have formed with a much higher eccentricity. A fraction of BBH mergers with such large eccentricities is only expected if they are driven by an environmental perturbation. Conversely, the presence of non-zero eccentricity at the minimum sensitive frequency of ground-based detectors is wholly inconsistent with the expected evolution of isolated massive binary stars (e.g., Belczynski et al., 2002). Due to these considerations, the detection of even a few sources with non-zero eccentricity (or the lack thereof; Zevin et al. 2021b) would act as a “tip of the iceberg” and could be used to put constraints on the fraction of mergers formed from dynamical channels in the entire observed population (cf., Baibhav et al., 2020).
@@
Tentative evidence for compact object mergers with non-zero eccentricity has recently been found in a handful of BBH merger events (e.g., Romero-Shaw et al., 2022; Gupte et al., 2024a; Iglesias et al., 2024; de Lluc Planas et al., 2025), though degeneracies with misaligned spin are, at present, hard to eradicate (e.g., Romero-Shaw et al., 2023) (although, see e.g. Divyajyoti et al. 2024; Morras et al. 2025). If this evidence becomes more robust, it would suggest that a somewhat large fraction of the BBH mergers detected thus far formed through a dynamical formation channel. Here, we consider a scenario in which additional eccentric BBH merger events are confidently detected, and investigate whether measuring their spin properties would further constrain which of the channels that produce eccentric mergers contribute. We focus on the spin properties because, for the subset of eccentric mergers, different formation channels make very distinct predictions. The two dimensionless spin vectors $\boldsymbol{\chi}_{1}$ and $\boldsymbol{\chi}_{2}$ of the black holes and their orbital angular momentum vector $\mathbf{L}$ are either expected to point in three random directions, or the spins are mutually closely aligned ($\boldsymbol{\hat{\chi}}_{1}\approx\boldsymbol{\hat{\chi}}_{2}$) but randomly oriented with respect to $\mathbf{\hat{L}}$, as argued below and illustrated in the right panel of Fig. 1.
@@
Figure 1: Left panel: Distribution of $\chi_{\rm eff}$ for mutually aligned spin directions (maroon) and random spin directions (teal). Dotted lines indicate equal-mass ($q=1$) and maximally spinning BBHs ($\chi_{1}=\chi_{2}=1$), solid lines assume an extended mass ratio distribution and a spin magnitude distribution that peaks at lower values, which are consistent with previous GW detections (see text). Black markers show three eccentric BBH merger candidates identified by Gupte et al. (2024a) (from bottom to top: GW200129, GW190701, and GW200208_22) and orange markers show four candidates identified by Romero-Shaw et al. (2022) (from bottom to top: GW200208_22, GW190521, GW190620, and GW191109). The commonly identified candidate GW200208_22 is highlighted by a filled marker (Romero-Shaw et al., 2025). Right panel: Sketch of uncorrelated (top) and mutually aligned black-hole spins of eccentric mergers (bottom).
@@
    Three random directions of $\boldsymbol{\chi}_{1}$, $\boldsymbol{\chi}_{2}$, and $\mathbf{L}$ are a natural property of eccentric BBH mergers that form in dense star clusters. There, BBHs that retain measurable eccentricity in the bandwidth of current GW detectors merge during resonant three- and four-body encounters that typically involve chaotic exchanges between binary members and scatterings that heavily perturb their orbits (Samsing et al., 2014; Antonini and Rasio, 2016; Rodriguez et al., 2018; Zevin et al., 2021b; Chattopadhyay et al., 2023). In these chaotic encounters, merging BBHs are randomly assembled without any correlations between the spins and orbital angular momentum. As a result, the directions of $\boldsymbol{\chi}_{1}$, $\boldsymbol{\chi}_{2}$, and $\mathbf{L}$ are independent of one another and effectively randomised.
@@
    In contrast, closely aligned spins $\boldsymbol{\hat{\chi}}_{1}\approx\boldsymbol{\hat{\chi}}_{2}$ and a random direction of $\mathbf{{L}}$ are expected for eccentric BBH mergers that result from the evolution of hierarchical triples. In these systems, the BBHs are formed from massive binary stars whose spins are likely (closely) aligned to one another by stellar processes. The gravitational perturbation from a tertiary companion promotes a subsequent merger by causing large-amplitude von Zeipel-Kozai-Lidov (ZKL) (Zeipel, 1909; Kozai, 1962; Lidov, 1962) oscillations of the BBH eccentricity and the orientation of its orbital angular momentum. BBHs that merge with a residual eccentricity undergo exceptionally strong ZKL oscillations while the spin evolution is highly non-adiabatic, i.e., the orbital angular momentum vector oscillates several orders of magnitude quicker than the spins could follow it through relativistic spin-orbit coupling (Antonini et al., 2018; Rodriguez and Antonini, 2018; Liu and Lai, 2018; Liu et al., 2019). As a result, the eccentric BBH mergers retain the mutually aligned spins they inherited from their binary progenitor stars while the direction of their orbital angular momentum vector is effectively randomised.
@@
In AGN disks, the spins of BBHs are expected to be mutually aligned through gas accretion due to the Bardeen-Petterson effect (Bardeen and Petterson, 1975), in which the alignment of the black-hole spin with a circum-black-hole disk proceeds much faster than the evolution of the spin magnitude (see McKernan et al. 2020; Santini et al. 2023 for related modeling). However, the directions of the orbital angular momentum can be misaligned with the black-hole spin directions depending on the frequency of binary-single interactions. This frequency is influenced by the properties of the AGN disks, especially the merging locations within the AGN disks. Recent models predict frequent interactions with random angular momentum directions (Tagawa et al., 2020a). In these cases, the spin-orbit alignment for the AGN channel is predicted to be similar to that of the three-body channel, and highly eccentric mergers are expected for a fraction of mergers. On the other hand, for the AGN channel, hierarchical mergers are inevitable (Yang et al., 2019; Tagawa et al., 2021; Gerosa and Fishbach, 2021), resulting in high masses and spins for these black holes, which can help distinguish them from triple systems.
In this work, we study how many confident detections of eccentric BBH mergers we would need before the two spin-orbit configurations of (i) three randomly distributed directions of $\boldsymbol{\chi}_{1}$, $\boldsymbol{\chi}_{2}$, and $\mathbf{L}$ and (ii) two randomly distributed directions of $\boldsymbol{\chi}_{1}$ and $\mathbf{L}$ and $\boldsymbol{\hat{\chi}}_{1}\approx\boldsymbol{\hat{\chi}}_{2}$ become statistically distinguishable. We highlight that focusing on spin-eccentricity correlations is particularly advantageous compared to previous approaches to distinguish formation channels from the marginalised spin distribution of all mergers (e.g., Rodriguez and Antonini, 2018; Rodriguez et al., 2019; Baibhav and Kalogera, 2024b). First, eccentric mergers can only be explained by a reduced number of formation channels, as any residual eccentricity close to merger is inconsistent with isolated binary formation scenarios (Fumagalli et al., 2024). Instead, eccentric mergers must have formed through a formation channel involving dynamical driving. Second, models of such formation channels only predict the clear distinction in spin distributions (i) and (ii) for eccentric events. The majority of events that merge on effectively circular orbits can originate from a wider variety of channels and have a less certain spin distribution. For instance, a significant fraction of all mergers that form in triples is expected to have a spin-orbit precession timescale comparable or shorter than the ZKL timescale, which leads to a more diverse merger distribution of spin-orbit angles (Liu and Lai, 2018; Rodriguez and Antonini, 2018). Meanwhile, an uncertain fraction of non-eccentric mergers in clusters may originate from primordial stellar binaries that experience no or only few dynamical interactions (Barber et al., 2024) and whose spin directions do not fully randomise as in (i). Thus, while any viable formation channel must be consistent with the entire joint spin-eccentricity distribution, this work shows that its detectably-eccentric sub-population is particularly insightful to constrain merger origins.
II Methods
@@
Random directions of $\mathbf{L}$, $\boldsymbol{\chi}_{1}$, and $\boldsymbol{\chi}_{2}$ correspond to uncorrelated distributions $\cos\theta_{1}\sim\mathcal{U}(-1,1)$ and $\cos\theta_{2}\sim\mathcal{U}(-1,1)$. In contrast, random directions of $\mathbf{L}$ but mutually aligned spins mean $\cos\theta_{1}=\cos\theta_{2}\sim\mathcal{U}(-1,1)$. We refer to these models as “uncorrelated” and “mutually aligned” respectively. The difference between the distribution of $\chi_{\rm eff}$ in both models is maximised if we consider the unrealistic but instructive case of equal-mass ($q=1$) and maximally spinning ($\chi_{1}=\chi_{2}=1$) BBHs. This is shown in the left panel of Fig. 1 (dotted lines), which illustrates that the mutually aligned model tends to produce more extreme values near $\chi_{\rm eff}\approx\pm 1$ whereas the uncorrelated model is more centred around zero. Both models resemble one another more closely if we consider more realistic, non-unity mass ratios and spin magnitudes. In Fig. 1, this is shown with solid lines where we draw the mass ratio from a power-law distribution $p(q)\sim q^{\alpha}$ between 0.1 and 1.0 and the spin magnitudes from some Beta distribution $p(\chi_{1,2})\sim\rm Beta(a,b)$. We choose fiducial values $\alpha=1.1$, $a=1.4$, and $b=3.6$ which are consistent with the distributions inferred from the GWTC-3 (Abbott and others, 2023). As a result, the mutually aligned model retains more pronounced tails of the distribution, whereas the uncorrelated model is more concentrated around $\chi_{\rm eff}=0$. Yet, the two distributions are more similar than in the previous case where $q=\chi_{1}=\chi_{2}=1$ due to a weaker contribution of the second term on the right-hand side of Eq. (1).
@@
For this purpose, we assume a total number of $N_{\rm tot}$ detected BBH mergers, which we set to $2,100$ as a reference for the 5th observing run of LVK (LVK O5) (Broekgaarden et al., 2024), $10^{4}$ for A#A\# (Abbott and others, 2020), and $10^{5}$ for the Einstein Telescope (Abac and others, 2025) and Cosmic Explorer (ET/CE) (Reitze and others, 2019). We assume that a fraction $0\leq\xi_{\rm ecc}\leq 1$ of them will retain confidently measurable eccentricity in the detector sensitivity window, resulting in $N_{\rm ecc}=N_{\rm tot}\times\xi_{\rm ecc}$ eccentric BBHs (rounded to the next integer). We only consider events in $N_{\rm ecc}$ with sufficient evidence for eccentricity, i.e., assume a suitable selection function was applied which discards events with marginal evidence to avoid spurious contamination by circular events (Mandel et al., 2019; Roy et al., 2025). At present, constructing a suitable selection function that would only collect confident detections of eccentricity is difficult given that current eccentric waveform models disagree on eccentricity in most events (Romero-Shaw et al., 2022; Gupte et al., 2024b). Here, we defer such task to future work and rely on the number $N_{\rm ecc}$ of confident eccentric detections that could be obtained in that way. For each detector configuration, we repeat the following steps $10^{4}$ times:
@@
    We produce a synthetic merger population by sampling $N_{\rm ecc}$ eccentric BBHs with mass ratios drawn from a power-law distribution $p(q)\sim q^{\alpha}$ between 0.1 and 1.0 and spin magnitudes independently drawn from some Beta distribution $p(\chi_{1,2})\sim\rm Beta(a,b)$. We assume that the distribution parameters are drawn from $\alpha\sim\mathcal{N}(1.1,1.0)$, $a\sim\mathcal{N}(1.4,0.1)$, and $b\sim\mathcal{N}(3.6,0.1)$. The assumed uncertainty is based on the current spin magnitude distribution of GWTC-3 (Abbott and others, 2023), but we note it will likely improve in the future as more events are being detected. For each of the eccentric BBH mergers we compute $\chi_{\rm eff}$ using Eq. (1) assuming either mutually aligned or uncorrelated spin directions. Lastly, we approximate the posterior distribution of each $\chi_{\rm eff}$ measurement with a truncated normal distribution between 1-1 and $1$, where the adopted standard deviation $\sigma_{\chi_{\rm eff}}$ represents typical measurement uncertainties.
@@
to quantify the confidence to recover the correct model, i.e., the mutually aligned model in the presence of mutually aligned data and the uncorrelated model in the presence of uncorrelated data. In the discussion below, we follow Kass and Raftery (1995) and consider $2\ln\mathcal{B}\gtrsim 6$ (i.e., $\log_{10}\mathcal{B}\gtrsim 1.3$) to indicate strong evidence for the correct model, $6\lesssim 2\ln\mathcal{B}\lesssim 6$ ($1.3\lesssim\log_{10}\mathcal{B}\lesssim 1.3$) no particular preference for either model, and $2\ln\mathcal{B}\lesssim-6$ ($\log_{10}\mathcal{B}\lesssim-1.3$)) strong evidence for the incorrect model.
@@
If not stated otherwise, we assume an eccentric merger fraction of $\xi_{\rm ecc}=0.05$ which agrees with predictions from dynamical channels for current ground-based detectors (e.g., Antonini et al., 2014; Zevin et al., 2021b), and is consistent with the order of magnitude of eccentric candidates in the current population of GW events (Romero-Shaw et al., 2022; Gupte et al., 2024a). However, it likely underestimates the fraction of eccentric mergers detectable with the next-generation detectors ET/CE, as they will be sensitive to smaller eccentricities (Dorozsmai et al., 2025). Below, we also study different values for $\xi_{\rm ecc}$. For the measurement uncertainty of $\chi_{\rm eff}$, we assume as a default $\sigma_{\chi_{\rm eff}}=0.2$ for LVK O5 and A# which is the typical uncertainty of the (quasi-circular) events identified in GWTC-3. We assume that ET/CE will allow measurements that are more precise by one order of magnitude, $\sigma_{\chi_{\rm eff}}=0.02$. Since it is unclear at what precision next-generation detectors will measure $\chi_{\rm eff}$} of eccentric events, we also study different values $\sigma_{\chi_{\rm eff}}$.
We note that step (i) implicitly assumes that eccentric and non-eccentric mergers share the same distribution of black-hole spin magnitudes, modelled as a Beta distribution. This assumption is justified because black-hole spin magnitudes are primarily determined by the physics of stellar core collapse, which is largely independent of the environment in which the binary forms. Different spin magnitudes may only arise from hierarchical mergers in dense stellar clusters, which can produce highly spinning black holes with $\chi_{i}\simeq 0.7$ (Gerosa and Fishbach, 2021). However, the contribution of such “second-generation” mergers to the eccentric subset as well as all mergers is expected to be small (e.g., Rodriguez et al., 2019; Mapelli et al., 2021).
It could also be possible that multiple formation channels contribute to the total merger rate. This would only significantly impact our analysis if these channels contributed comparably and had substantially different spin magnitude distributions. Such a scenario would require a degree of fine-tuning. In contrast, the eccentric merger fractions $\xi_{\rm ecc}$ considered in this work are consistent with all mergers being dominated by dynamical formation processes. In these channels, only a subset of binaries retains measurable eccentricity within the sensitivity range of our detectors, while the rest formed through the same mechanism but circularized before entering the detectable eccentricity regime. Moreover, any alternative formation channel would need to produce a markedly different spin distribution, which is unlikely given that spin magnitudes are predominantly governed by core-collapse physics.
III Results